# Sell It Before You Make It: Revolutionizing E-Commerce with Personalized AI-Generated Items

### v1

@@ A.3. Relation With DPO
When we only have one positive image and one negative image, i.e., $|\mathcal{P}|=|\mathcal{N}|=1$, the training objective in Eq. 18 can be reduced to the classical direct preference optimization (DPO) objective for diffusion models proposed by Wallace et al. (2024). The derivation is as follows:
@@ Appendix B Experiment Setups
B.1. Datasets
We adopt the public PickaPic dataset (Kirstain et al., 2023) and Alibaba’s Industrial dataset for experiments. We show the dataset statistics in Table 5.
Table 5. The dataset statistics.
Dataset #Users #Prompts #Samples Group Size
PickaPic 3,094 33,164 529,219 2
Industrial 54,995 60,616 607,328 5
B.2. Baseline Models
For user preference estimation task, we adopt Aesthetics (Schuhmann, 2022), CLIP-H (Radford et al., 2021), ImageReward (Xu et al., 2024a), HPSv1 (Wu et al., 2023b), HPSv2 (Wu et al., 2023a) and PickScore (Kirstain et al., 2023) as the baseline models:
@@ B.2. Baseline Models
For personalized image generation task, we adopt Stable Diffusion v1.5 (Rombach et al., 2022), SFT (Ho et al., 2020), DPO (Wallace et al., 2024), SPO (Liang et al., 2024) and PMG (Shen et al., 2024) as baselines:
@@ B.3. Evaluation Metrics
For the personalized image generation task, we adopt Aesthetic (Schuhmann, 2022), CLIP Score (Radford et al., 2021), PickScore (Kirstain et al., 2023), and PerFusionRM as our metrics.
@@
B.4. Implementation Details
For PerFusionRM on the user preference estimation task, we use PickScore (Kirstain et al., 2023) as the backbone model, which is an improved version of CLIP to better capture the general human preference. We adopt AdamW (Loshchilov, 2017) as the optimizer. The batch size is set to 128. The learning rate is set to 1e-3 with a 500-step warmup. The weight decay is set to 1e-2.
For the personalized image generation task, we first conduct direct preference optimization (DPO) (Wallace et al., 2024) on Stable Diffusion v1.5 to capture the general user preference, which is then adopted as the backbone model for our PerFusion. We train the model at fixed square resolutions. We adopt AdamW (Loshchilov, 2017) as the optimizer. We set the learning rate to 1e-8 with a 500-step warmup. The weight decay is set to 1e-2. The batch size is selected from $\{256,512,1024\}$. The hyperparameter $\beta$ is chosen from $\{1500,2000,2500,3000\}$.

### v2

@@ A.3. Relation With DPO
When we only have one positive image and one negative image, i.e., $|\mathcal{P}|=|\mathcal{N}|=1$, the training objective in Eq. 18 can be reduced to the classical direct preference optimization (DPO) objective for diffusion models proposed by wallace2024diffusion. The derivation is as follows:
@@ Appendix B Experiment Setups
B.1. Baseline Models
For user preference estimation task, we adopt Aesthetics (aesthetic_classifier), CLIP-H (radford2021learning), ImageReward (xu2024imagereward), HPSv1 (wu2023human), HPSv2 (wu2023humanv2) and PickScore (kirstain2023pick) as the baseline models:
@@ B.1. Baseline Models
For personalized image generation task, we adopt Stable Diffusion v1.5 (rombach2022high), SFT (ho2020denoising), DPO (wallace2024diffusion), SPO (liang2024step) and PMG (shen2024pmg) as baselines:
@@ B.2. Evaluation Metrics
For the personalized image generation task, we adopt Aesthetic (aesthetic_classifier), CLIP Score (radford2021learning), PickScore (kirstain2023pick), and PerFusionRM as our metrics.
@@
B.3. Implementation Details
For PerFusionRM on the user preference estimation task, we use PickScore (kirstain2023pick) as the backbone model, which is an improved version of CLIP to better capture the general human preference. We adopt AdamW (loshchilov2017decoupled) as the optimizer. The batch size is set to 128. The learning rate is set to 1e-3 with a 500-step warmup. The weight decay is set to 1e-2. PickScore and our PerfusionRM are tuned on both datasets, while other baselines stay untuned to capture different aspects of user vision preferences, e.g., aesthetics and text-image alignment.
For the personalized image generation task, we first conduct direct preference optimization (DPO) (wallace2024diffusion) on Stable Diffusion v1.5 to capture the general user preference, which is then adopted as the backbone model for our PerFusion. We train the model at fixed square resolutions. We adopt AdamW (loshchilov2017decoupled) as the optimizer. We set the learning rate to 1e-8 with a 500-step warmup. The weight decay is set to 1e-2. The batch size is selected from $\{256,512,1024\}$. The hyperparameter $\beta$ is chosen from $\{1500,2000,2500,3000\}$. All baseline models use the same input prompts for fair comparison, except PMG, which is a method to improve input prompts based on user profiles.